OpenStudy (anonymous):

-2/3(4x+5)= -10/3 Is this correct? ugh I hate fractions

OpenStudy (anonymous):

not sure what "is this correct" means it is an equation, we can solve it if you like

OpenStudy (anonymous):

haha yeah it means solve please because i'm just checking my work

OpenStudy (anonymous):

if it is \[-\frac{2}{3}(4x+5)=-\frac{10}{3}\] then we can get rid of the annoying fractions (which everyone hates) by multiplying both sides by \(-3\) and starting with \[2(4x+5)=10\]

OpenStudy (tkhunny):

It is absolutely not correct. Hatred is an ugly thing. Let it go. Embrace the fractions,

OpenStudy (anonymous):

then it should be real easy can almost do it in your head

OpenStudy (anonymous):

hahaha sorry @tkhunny and thanks @satellite73

OpenStudy (anonymous):

divide both sides by \(2\) and get \(4x+5=5\) which can only be the case if \(x=0\)